In Solution of Traingle,
How doess a*cos2(C/2) + c*cos2(A/2)= 3b/2
become a(1+cosC) + c(1+cosA)=3b?

Grade:12

1 Answers

Rinkoo Gupta
9 years ago
acos2(C/2)+c
cos2(A/2)=3b/2
2acos2(C/2)+2ccos2(A/2)=3b
a(1+cosC)+c(1+cosA)=3b
Using formula
2cos2x=1+cos2x

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty